OpenStudy (briensmarandache):

Partial fraction decomposition (2x^2-x+4)/(x^3+4x) I got a=1 b=1 c=-1 which is wrong I'm sure

OpenStudy (mathmath333):

its correct i think

OpenStudy (mathmath333):

\(\large\tt \color{black}{\dfrac{2x^2-x+4}{x^3+4x}=\dfrac{1}{x}+\dfrac{x-1}{x^2+4}}\)

OpenStudy (briensmarandache):

Ok so I'm glad I got that portion of the problem right lol thanks